# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ 3.3 Examples and Further properties of the (Fleischer) Net Riemann Integral
Further examples of the Net Riemann integral, such as the Riemann-Stieltjes integrals of Candeloro [2], Bochner and Ky-Fan [78], and the abstract Riemann integral of Mcshane [54], can be specified and show to be contained in our main integral concept. We leave further details for another work. Some integrals that don’t seem to fit in the general Riemmanian integration schemes presented until now will be pointed out in the last sections of this work.
As a way of obtaining another useful and interesting example, at least in a restricted sense, we now formulate an equivalence result concerning the Net Riemann integral and the $S^{*}$-integral. This aforementioned equivalence is of a restricted nature because it shows that the existence of the $S^{*}$-integral implies the existence of a particular example of the Net Riemann integral when $Z$ equipped with a Boccuto-Candeloro net convergence structure which is topological. In fact, we have:
@@
Let $f\in L^{1}_{S^{*}}(\mu)$ be a $S^{*}$-integrable function in the context of conditions (i) and (ii) of Definition 3.15, where $Z$ is equipped with a Boccuto-Candeloro net convergence structure which is generated by a vector space topology. Then, $f$ is Sion integrable.
@@
As the convergence in $Z$ is topological, we may find a topology $\tau$ for which each convergence net in $Z$ is convergent in this vector topology. Denote by $\mathcal{U}$ the filter of neighborhoods of $\tau$. In this case, as $f$ is $S^{*}$-integrable in $\Omega$ (the proof for other sets is immediate), fix an arbitrary $U\in\mathcal{U}$ and take symmetric and absorbent $V\in\mathcal{U}$ such that $V+V\subseteq U$ and $P_{0}\in\mathcal{P}(\Omega)$ such that, denoting by $s=\int_{\Omega}^{S^{*}}fd\mu$,
$\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-s\in V,$
@@
$\sum_{\alpha\in\Delta(P)}f(\delta_{\alpha})\mu(\alpha)\in s+U.$
@@
$\int_{\Omega}^{S^{*}}fd\mu=s=\int_{\Omega}^{(Si)}fd\mu$
It follows that problem (iv) in Observation 3.2 is solved in the topological case, and we get a new example of the Net Riemann integral with $Z$ equipped with a topological convergence. In particular, we obtain as a "methodological corollary" that the Net Riemann integral can be used effectively in Hewitt-Yosida, Lebesgue and other decompositions for Riesz Spaces - valued set functions, as in [79].
It would be interesting to obtain an answer to the following question regarding the preceding discussion: for which Boccuto-Candeloro structures, besides the topological ones, is Proposition 3.12 true ?
We end this section with some comments on the behavior of the Net Riemann integral on hereditary (i.e, subset) and simple functions integration.
First, we notice that, as indicated before, and commented by Szabó and Száz [22], the usual integration formula for simple functions in point (ii) of Observation 3.2 will not hold for the Net Riemann integral. That can be seen, for example, from the fact that it contains the Henstock-Kurzweil integral in the case of Riesz valued functions and topological scalar measures (see [11]), and even in this case it is necessary to suppose the Radon regularity of the measure to obtain such property. This motivates the next concept, taken from Section 4 of [22]:
Definition 3.17 (Regular Integrators).
$\int_{\Omega}\mathbbm{1}_{A}d\mu=\mu(A),$
for each $A\in\mathcal{H}$.
It follows directly from (i) of Theorem 3.1 and the properties of the bilinear product in Definition 3.6 that simple functions are then Net Riemann Integrable, that is:
Proposition 3.13 (Integration of Simple Functions).
Let $f:\Omega\rightarrow X$ be a $\mathcal{H}$-simple function, i.e:
$f=\sum_{i=1}^{n}x_{i}\mathbbm{1}_{A_{i}},$
$\int_{\Omega}fd\mu=\sum_{i=1}^{n}x_{i}\mu(A)$
Lastly, we have the following question of hereditary integration relating to the main integral, i.e: if a function $f$ is Net Riemann integrable on $B\in\mathcal{H}$ and $A\in\mathcal{H}$ is such that $A\subseteq B$, does it implies that $f$ is integrable on $A$ ? As the example of the Riemann-Stieltjes integral of [54] shows, this cannot be positively answered in general.
If $\Gamma$ is, for example, the space of countable coverings/partitions in the case of Kolmogorov’s $S^{*}$-integral, this property can be proved in full generality of net convergence structures, as seen by Theorem 3 of [14]. Similarly for Theorem 53 of [60] on Page 32 and Theorem 2.2 of [75].
At this point it is ideal to comment further the points made in Observation 3.1. For the Fleischer Net Riemann integral, the hereditary integration property, by which we call the existence of integrals for subsets which are contained in integrable ones, seems difficult to obtain without imposing more conditions on the process of associating a Fleischer-Szabó-Száz net to each set $A\in\mathcal{H}$. For specific integrals, such as the ones presented in Examples 3.2-3.6, this result is valid (see the last paragraph).
For our case of general Boccuto-Candeloro net convergence structures and the Fleischer Net Riemann integral, the analogue result of 5°paragraph of [24] seems the best we can get at the moment, which is, in some sense, a result of imposing a Cauchy type condition for the Riemann sums.
Theorem 3.2 (Subset Integration - Fleischer Net Riemann Integral).
Suppose that $Z$ is complete in its net convergence structure. In the context of Fleischer’s modification of Definition 3.9 of the Net Riemann integral in a fixed partially ordered convergence structure, fix $A,B\in\mathcal{H}$ such that $A\subseteq B$. Consider the following hypothesis:
@@
    For each $\gamma_{B},\gamma_{B}^{{}^{\prime}}\in\Gamma(B)$, there exists $\gamma_{A},\gamma^{{}^{\prime}}_{A}\in\Gamma(A)$ and $\gamma_{A^{C}}\in\Gamma(A^{c})$ such that there exists $\gamma^{1}_{B},\gamma^{2}_{B}\in\Gamma(B)$ such that $\gamma^{1}_{B}\geq\gamma_{B},\gamma^{2}_{B}\geq\gamma^{{}^{\prime}}_{B}$ with corresponding Fleischer-Szabó-Száz nets
    $\gamma^{1}_{B}\rightarrow((\sigma_{\gamma_{A}}\cup\sigma_{\gamma_{A^{c}}}),(\tau_{\gamma_{A}}\cup\tau_{\gamma_{A^{c}}})),$
    $\gamma^{2}_{B}\rightarrow((\sigma_{\gamma^{{}^{\prime}}_{A}}\cup\sigma_{\gamma_{A^{c}}}),(\tau_{\gamma^{{}^{\prime}}_{A}}\cup\tau_{\gamma_{A^{c}}})),$
Then, if $f$ is Fleischer Net Riemann integrable on $B$, it is Net Riemann Integrable on $A$ and:
$\int_{B}fd\mu=\int_{A}fd\mu+\int_{A^{c}}fd\mu$
The strategy of the proof being the same as the already cited results of Theorem 5.3 of [60] on Page 32 and Theorem 2.2 of [75], which are applications of Cauchy nets, we omit it. When a topology or order-type convergence is available, weaker conditions for the validity of the theorem can de found - besides the references above, an archetypic reference would by Proposition 5.2.10 of [10]. Finally, we point out that this result can also be modified to obtain additivity of the set function induced by the Fleischer Net Riemann integral and the Net Riemann integral, as in the analogous case of references [24] and [75], but we will not comment further.
Having finished the results about the fundamental (or "basic") properties of the Net Riemann integral, we now discuss some convergence theorems for the (Fleischer) Net Riemann integral in the context of specific partially ordered net convergence structures.
4 Uniform Convergence Theorems for Net Riemann and $S^{*}$-partition Integrals
To obtain more refined properties of the (Fleischer) Net Riemann integral and correlated concepts introduced in previous sections, we need some convergence theorems, specially (order) uniform ones. The purpose of these convergence theorems are twofold:

### v2

@@ Abstract
Further examples of the Net Riemann integral can be constructed readily by the same methods above. Indeed, we shall enumerate only those integration procedures that can be fitted, in a method of construction, in some of the examples above. We also stress that, in Example 2.2, we noticed the fact that the net Riemann integral contain the Lebesgue type integrals of Phillips (1940) and Rickart (1942). Some more examples of the Lebesgue case will be given later.
  1. 1.
    The Riemann-Stieltjes, Moore-Pollard, Young and Young-Stieltjes type integrals of Candeloro (1996), Hayes (1974), Gowurin (1936) and Dudley and Norvaiša (2011), as well as the finite partition refinement integral of Edwards and Wayment (1974) (which they call the $v$-integral).
  2. 2.
    The multilinear integrals of Young, Stieltjes and Moore-Pollard in the sense of Halilovic (1990), developed further in Halilović (1991), and more recently in Halilović (2009).
  3. 3.
    The Burkill-Cesari-Henstock integrals of Henstock (2017), Sworowski (2006) Boccuto and Sambucini (1996) and Skvortsov et al. (2013). In fact, Chernyavskij (1975) and Chernyavskiĭ (1965) obtains a result that the Kolmogorov $S$-integral (i.e, for finite partitions) contains the group-valued integral of Burkill.
  4. 4.
    The finite partition refinement type integral of Saeks and Goldstein (1972), called the Cauchy integral.
  5. 5.
    The (SL), metric-semigroup valued and topological Henstock-Kurzweil integrals of Boccuto and Riečan (2006b), Boccuto et al. (2009), Boccuto and Riečan (2006c), Boccuto and Riečan (2006a) and Vrábel and Vrábelová (2011) (omitting some more publications of the same authors, but with the same integrals).
  6. 6.
    The Saks type integral of Bochner and Fan (1947).
  7. 7.
    The integral of Robdera (2025), which is based on nets of partitions (i.e $\Gamma$ is a directed set based on partitions).
  8. 8.
    The Riemann partition integral of Erben and Grimeisen presented in Erben (1981), Erben and Grimeisen (1990), Erben and Grimeisen (1991).
  9. 9.
    The general Riemann integral of Spaltenstein (1995), which covers the Denjoy integral and other generalized (Riemann) integrals.
  10. 10.
    The BV-sets and geometric theory of Riemann integration of Pfeffer (1993), as well as the locally compact Hausdorff space integrals of Ahmed and Pfeffer (1986), Pfeffer (1969) and Pfeffer (1970).
  11. 11.
    The non-tagged gauge integral of McShane (1969) and McGill (1980), as well as the Mcshane integral in quasi-Radon measure spaces of Rodríguez (2006), Fremlin and Mendoza (1994), Reynolds Jr (1997), Fremlin (1995) and Fremlin (1994) with a textbook treatment of the Banach case contained in Schwabik and Ye (2005). Also, the $H$-integral of Ridder, which is strictly general than the Henstock-Kurzweil integral, is contained in our integral by the results of Tack-Wang (1972).
  12. 12.
    The totalization integral of Trjitzinsky (1969).
  13. 13.
    The topological semigroup-valued truncation integrals of Goguadze (1983), Goguadze (1984), Impens (1978) and Delanghe and Impens (1979), which are constructed in the same way as in Sion (1973).
  14. 14.
    The product Kolmogorov integrals of Chernyavskij (1961), Areshkin and Koroleva (1983) and Ferreira (2004) in the form of the abstract Riemann net integral for (commutative) products in groups.
Therefore, the net Riemann integral provides a uniquely and very general integration procedure in integration theory literature, which can be used in its various specializations to solve diverse problems in analysis (as can be gauged from the references just cited). A cursory look at the vast bibliography and examples of Henstock (1991) will also produce many more examples, as his division space integral is included in ours.
We finish this section by exposing two shortcomings of these examples:
  1. (i)
    The relation between them and the Lebesgue, and other types of non-Riemann integration procedures, has not been fully clarified.
  2. (ii)
    There are integrals, such as Kolmogoroff (1930) type $II$ integral, Birkhoff (1935) and Goguadze (1979) integrals, the integrals of Popescu (1970a), Popescu (1963), Popescu (1965), Popescu (1966), Popescu (1970b) and Ionescu Tulcea (1954b), Ionescu Tulcea (1954a), Ionescu Tulcea (1953) and Deleanu (1957) that use infinite Riemann sums in the approximating scheme, and therefore are not contained ("directly") in the examples above and haven’t been studied much in the literature.
We study first point (ii), and then prove convergence theorems for the net Riemann integral and the infinite sums Riemann type-integrals to obtain a possible solution to (i).
3 Summability in Riesz Spaces and Integrals Based on Infinite Riemann Sums.
In this section, we shall study a type of integral that uses infinite Riemann sums to base its approximation. Various integrals in the literature use this approach (see point (ii) in the ending of last section), but only is some restricted settings such as locally convex spaces (see Popescu (1970a)) or real functions and measures (see Goguadze (1979) for an in depth study of this case). To obtain such integration procedure in our case, we have to first develop the rudiments of the concept of conditionally and unconditionally summable series, as well as their properties, in convergence spaces. Besides the references cited in this section, we draw the reader attention to the concept of abstract axiomatic sums in Freni (2023), Nielsen (2025), Katz and Straus (1965) and Andrés-Martínez and Heunen (2025).
To begin the preparations to the integral, we note that, as said above, we have two notions of converging series, one corresponding to unconditional convergence, and other to conditional convergence, as in the classical reference of Bourbaki (2013). In this context, consider the next definition:
Definition 3.1 (Unconditionally Convergent Series).
Let $(X,\eta)$ be a vector space equipped with a net convergent structure $\eta$. Then, we say that a net $\{x_{i}\}_{i\in I}$ in $X$, with partial (unordered) sums of elements (in $X$)
$\{\sum_{j\in J}x_{j}\}_{J\in\mathcal{F}(I)},$
$x=\lim_{(\mathcal{F}(I),\subseteq,\eta)}\sum_{j\in J}x_{j}$
Definition 3.2 (Conditionally Convergent Series).
Let $(X,\eta)$ be a vector space equipped with a net convergent structure $\eta$. Then, we say that a sequence of elements $\{x_{n}\}_{n\in\mathbb{N}}$, with partial (ordered) sums of elements (in $X$)
$\{\sum_{i=1}^{n}x_{i}\},$
$x=\lim_{(\mathbb{N},\eta)}\sum_{i=1}^{n}x_{i}$
Similar notions, specifically the first one, can be found in Higgs (1988), Brunker (1980) and Hebisch (1992). We now prove some connections between these two concepts in general net convergence structures.
Let $(X,\eta)$ be a vector space equipped with a net convergent structure $\eta$. and $\{x_{n}\}_{n\in\mathbb{N}}$ a sequence with unconditionally convergent series. Then, it has a conditionally convergent series with same limit.
Proof.
We prove this result using a quasi-subnet argument. In fact, suppose that $\{x_{n}\}_{n\in\mathbb{N}}$ has a unconditionally convergent series, and take an arbitrary $J\in\mathcal{F}(\mathbb{N})$. Choose $N_{0}=\{1,\cdots,\max(J)\}$. Then, if we take $n\in\mathbb{N}$ with $n\geq\max(J)$, it is clear that:
$\{\sum_{i=1}^{n}x_{i}\}_{n\geq\max(J)}\subseteq\{\sum_{k\in K}x_{k}\}_{J\subseteq K},$
which shows that $\{\sum_{i=1}^{n}x_{i}\}_{n\in\mathbb{N}}$ is a quasi-subnet of $\{\sum_{j\in J}x_{j}\}_{J\in\mathcal{F}(\mathbb{N})}$. As this last one is convergent to $x\in X$, it follows that, by axiom (ii) of Definition 1.13, the sequence has a conditionally convergent series to $x$. ∎
Regarding to the tail of a series, we have, by a similar argument:
Proposition 3.2 (Tail of Series).
Let $(X,\eta)$ be a vector space equipped with a net convergent structure $\eta$. and $\{x_{n}\}_{n\in\mathbb{N}}$ a sequence with a conditionally convergent series. Then, the tail of the series, i.e., the sequence:
$\{\sum_{i=n}^{\infty}x_{i}\}_{n\in\mathbb{N}},$
is convergent (with respect to $\eta$) to 0.
Both results, specially Proposition 3.1, are interesting in the context of measure theory in partially ordered spaces with net convergence structures compatible with order (in the sense of the last section). Regarding this context, the next definition is also useful:
Definition 3.3 (Monotonically Convergence Property).
Let $(X,\eta)$ be a Riesz space equipped with a partially ordered net convergent structure $\eta$. Then, we say that $(X,\eta)$ has the monotonically convergence property (MCP) if, given an increasing bounded above net with a supremum $\{x_{i}\}_{i\in I}$, it converges to its supremum in $\eta$, i.e.
$\lim_{(I,\eta)}x_{i}=\sup_{i\in I}x_{i}$
This is motivated by the following classical result, which can be found in various references, such as Pavlakos (1983), Papangelou (1964), Luxemburg and Zaanen (1971), Boccuto and Dimitriou (2015), Boccuto et al. (2012b) (for further examples, see specially Boccuto and Dimitriou (2015)):
Proposition 3.3 (Spaces with the MCP).
If $X$ is a Dedekind complete Riesz space with partially ordered net convergence structure $\eta$ being $(o_{3})$ or $(o)$ convergence, then it has the MCP property.
We also have the following example of a topological convergence with the MCP, showing that the concept is not restricted to spaces with a "purely" order based convergence.
Example 3.1.
Let $(X,\eta)$ be the space $L^{0}(\Omega,\Sigma,\mathbb{P})$ of (equivalence classes of) real Borel-measurable functions defined on a probability space, equipped with convergence in probability, and the partial ordering of almost sure pointwise inequality. It is not difficult to check that this convergence is topological and gives a partially ordered net convergence structure in $X=L^{0}(\Omega,\Sigma,\mathbb{P})$. In this case, taking $\{X_{n}\}_{n\in\mathbb{N}}$ a monotone sequence of random variables in $X$ bounded above (by a random variable $Y$), it follows from Theorem 3.5 of Gut (2013) that this sequence has a limit in probability given by its supremum. It follows that the present $(X,\eta)$ has the MCP.
In this case, as we will use repeatedly, we have the following result, which can be proved as in Lemma 1.2 of Pavlakos (1983) using a quasi-subnet argument:
Proposition 3.4 (Non-negative Series Convergence).
Let $(X,\eta)$ be a Riesz space equipped with a (locally solid) net convergent structure $\eta$ with MCP property and $\{x_{n}\}_{n\in\mathbb{N}}$ a non-negative sequence with conditionally convergent series and partial sums $\{\sum_{i=1}^{n}x_{i}\}_{n\in\mathbb{N}}$ bounded above in $X$. Then, the series associated to $\{x_{n}\}_{n\in\mathbb{N}}$ is unconditionally convergent and:
$\sum_{i=1}^{\infty}x_{i}=\lim_{(\mathcal{F}(\mathbb{N}),\subseteq,\eta)}\sum_{j\in J}x_{j}=\sup_{n\in\mathbb{N}}\sum_{j=1}^{n}x_{i}$
Observation 3.1 (Ideal Infinite Element).
We can dispense with the existence of the supremum, or the MCP property, in the results above by annexing an ideal element $\infty$ to $X$ in such a way that every increasing sequence has a supremum equal to it. Another alternative would be to work with the Dedekind completion of such space (see Fleischer (1988) for a general treatment of this case).
@@
Let $(X,\eta)$ be a Dedekind complete Riesz space where $\eta$ is the partially ordered net convergence structure given by $(D)$-convergence. Then, a non-negative sequence $\{x_{n}\}_{n\in\mathbb{N}}$ has a conditionally convergent series if and only if it has an unconditionally convergent series with the same limit.
@@
$\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}x_{i}-\sum_{i=1}^{n}x_{i}\leq p_{n},$
$\sum_{j\in J_{0}}x_{j}=\sum_{i=1}^{k}x_{i}\leq\sum_{j\in J}x_{j}.$
Therefore,
These results allow us to construct a theory of $\sigma$-additive measures with values in spaces with a general net convergence structure, and having further interesting properties in presence of a partial order. In this regard, we now state a fundamental definition:
Definition 3.4 ($\sigma$-additive Measure).
Let $(\Omega,\mathcal{H})$ a paved space and $(Y,\eta)$ be a set equipped with a net convergence structure $\eta$. Then, we say that a set function:
$\mu:\mathcal{H}\rightarrow Y,$
Its important to notice that the series in the definition above is a unconditional one, because the use of the (commutative operation of) countable union in its definition - a point forcefully put in the topological context by Fleischer and Traynor (1989), and which gives interesting consequences in case we consider measure theory in non-commutative structures (see Pap (1995)).
By using Propositions 3.1 and 3.2, its possible to prove the following fact directly (as it depends only in the series definition):
Proposition 3.6 (Monotone Continuity and $\sigma$-Additivity).
Let $(\Omega,\mathcal{H})$ a paved space by a ring and $(Y,\eta)$ be a set with partially ordered net convergence structure $\eta$. Then, a finitely additive set function:
$\mu:\mathcal{H}\rightarrow Y^{+},$
$\lim_{(\mathbb{N},\eta)}\mu(A_{n})=0$
If $Y$ is a Riesz space equipped with a partially ordered net convergence structure and $\mu$ assumes values in the non-negative cone of $Y$, the limit is decreasing in $n$.
It follows that our definition of $\sigma$-additive measure is the same as the one used in the theory of non-negative $\sigma$-additive measures in Riesz spaces equipped with order or (D)-convergence - see Section 2.3 of (Boccuto and Dimitriou, 2015, Chapter 2) for details.
Having a notion of series in spaces with abstract convergence structures allow us to define a different type of integral that is different in nature compared to the Net Riemann integral.
Before the definition itself, consider the following auxiliary object and convention:
Definition 3.5 (Choice Function).
Also, for a countable collection of sets $P=\{\alpha\}_{\alpha\in P}$, we denote for $J\in\mathcal{F}(\mathbb{N})$,
$\sum_{j\in J,\alpha\in P}g(\alpha_{j}),$
where $g$ is a set function, to be the sum of the elements of $P$ indexed by $J$.
Now we may define a new integral, which in its original form is due to Kolmogoroff (1930), but the modern usage of it is due (initially) to Deleanu (1957), and Ionescu-Tulcea in Ionescu Tulcea (1954b) and Ionescu Tulcea (1954a), called the $S^{*}$-integral:
Definition 3.6 (Kolmogorov $S^{*}$ integral).
Given and ordered pair:
consisting of
  1. (i)
    $(\Omega,\mathcal{H})$, a paved space $\mathcal{H}$;
  2. (ii)
    $(X,\eta_{1}),(Z,\eta_{2})(Z,\eta_{3})$, spaces with (partially ordered, if Riesz spaces) net convergence structures connected by a bilinear product $\bullet$;
The integral in Definition 3.6 may be paraphrased as follows:
Definition 3.7 ($S^{*}$-integral Unconditional Series).
In other words, whenever the $S^{*}$-integral exists, we assume the existence of the series on the integral for partitions at least cofinal for the convergence (i.e, which refines a fixed element of $\mathcal{P}(\Omega)$). We also comment that its possible to generalize the partitions above to specific classes generated by other means (i.e, gauges and similar techniques) and obtain integrals in the sense of Dubois and M’Khalfi (1988b) and M’Khalfi (1989).
For the rest of the text, we denote by $L_{S^{*}}(\mu)$ the space of all functions $f:\Omega\rightarrow X$ that are $S^{*}$-integrable with respect to the (fixed) set function $\mu:\Omega\rightarrow Y$ on each $A\in\mathcal{H}$ with the integration structures of Definition 3.6.
We now summarize some properties of the $S^{*}$-integral in this abstract setting. As the proofs are exactly the same as in the net Riemann integral, we omit them. Some of these may be found (in a different and less general form) in Deleanu (1957).
Proposition 3.7 (Properties of the $S^{*}$-Integral).
The $S^{*}$-Integral has the following properties, where we assume $X,Y,Z$ Riesz spaces in items with inequalities:
  1. (i)
    Let $f,g\in L_{S^{*}}(\mu)$. Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}(f+g)d\mu=\int^{S{*}}_{A}fd\mu+\int^{S{*}}_{A}gd\mu,$
    i.e, the integral is additive on the space $L_{S^{*}}(\mu)$;
  2. (ii)
    Let $f\in L_{S^{*}}(\mu)$ such that $f\geq 0$ and $\mu\geq 0$. Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}fd\mu\geq 0,$
  3. (iii)
    Let $\mu\geq 0$ and $f,g\in L_{S^{*}}(\mu)$ such that:
    $g\geq f.$
    Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}gd\mu\geq\int^{S^{*}}_{A}fd\mu,$
    i.e, the integral is an isotone map on $L_{S^{*}}(\mu)$ equipped with the partial ordering inherited from $X$.
  4. (iv)
    Let $\mu\geq 0$ and $f\in L_{S^{*}}(\mu)$ such that $|f|\in L_{S^{*}}(\mu)$, then for each fixed $A\in\mathcal{H}$:
    $|\int^{S^{*}}_{A}fd\mu|\leq\int^{S^{*}}_{A}|f|d\mu.$
  5. (v)
    $\int^{S^{*}}_{A\cup B}fd\mu=\int^{S^{*}}_{A}fd\mu+\int^{S^{*}}_{B}fd\mu$
  6. (vi)
  7. (vii)
    $\int_{A}^{S^{*}}fd\mu=c\mu(A)$
What is interesting about this integral is that it gives $\sigma$-additivity automatically in the presence of a summability condition, which Deleanu (1957) (in section $4$ of the article, for which we refer the reader for details) calls the ()(*)-condition, and is satisfied for any complete (Hausdorff) topological group or/and Dedekind complete lattice ordered group. In fact, he obtain the following result:
Theorem 3.1 ($\sigma$-additivity of the $S^{*}$-integral).
Let $f:\Omega\rightarrow X$ be a $S^{*}$-integrable function with respect to a(n) (arbitrary) set function $\mu:\mathcal{H}\rightarrow Y$, where $\mathcal{H}$ is a $\sigma$-ring of subsets of $\Omega$ and the integrability occurs with respect to a fixed net convergence structure where $Z$ satisfies the ()(*)-condition. Then, the set function defined on $\mathcal{H}$:
$\nu^{S^{*}}_{f}({A})=\int_{A}^{S^{*}}fd\mu,$
is a $\sigma$-additive measure with values in $Z$.
We refer the reader to (Deleanu, 1957, Page 33) for a proof.
We now formulate a result to connect the net Riemann integral and the $S^{*}$-integral. This aforementioned equivalence is of a restricted nature because it shows that the existence of the $S^{*}$-integral implies the existence of a particular example of the Net Riemann integral when $Z$ equipped with a net convergence structure that is topological (notice that $X$ and $Y$ need not be of this type). In fact, we have the following result that proves a fact cited by Sion (1973), on Page 16, without proof:
Proposition 3.8.
Let $f\in L_{S^{*}}(\mu)$ be a $S^{*}$-integrable function in the context of conditions (i) and (ii) of Definition 3.6, where $Z$ is equipped with a net convergence structure which is generated by a vector space topology. Then, $f$ is Sion integrable [^13] Here and thereafter, we denote the Sion integral of $f$ on $\Omega$ by $\int_{\Omega}^{(Si)}fd\mu$.
Proof.
@@
As the convergence in $Z$ is topological, we may find a topology $\tau$ for which each convergence net in $Z$ is convergent in this vector topology. Denote by $\mathcal{U}$ the filter of absorbent and symmetric neighborhoods of $\tau$. In this case, as $f$ is $S^{*}$-integrable on $\Omega$ (the proof for other sets is immediate), fix an arbitrary $U\in\mathcal{U}$ and take symmetric and absorbent $V\in\mathcal{U}$ such that $V+V\subseteq U$ and $P_{0}\in\mathcal{P}(\Omega)$ such that, denoting by $z=\int_{\Omega}^{S^{*}}fd\mu$,
$\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-z\in V,$
@@
$\sum_{\alpha\in\Delta(P)}f(\delta_{\alpha})\mu(\alpha)\in z+U.$
@@
$\int_{\Omega}^{S^{*}}fd\mu=z=\int_{\Omega}^{(Si)}fd\mu$
In the partially ordered case, we can prove a similar statement, but with a further restriction. To motivate this restriction, we note that several limit theorems for lattice-group/Riesz space-valued measures can be proved only if we choose a $(D)$-sequence for convergence that is uniform in some sense: either the same for a family of measures or for a collection of sets. This is, for example, the technique of Boccuto and Candeloro (2002) to prove a type of Vitali-Hahn-Saks theorem for lattice group-valued measures, the same occurs in Boccuto et al. (2012a). Intuitively, we need this to control simultaneously an uncountable number of nets, and reduce the case to an application of Fremlin’s lemma. In this setting, then, we have:
Proposition 3.9.
Let $f\in L_{S^{*}}(\mu)$ be a $S^{*}$-integrable function in the context of conditions (i) and (ii) of Definition 3.6, where $Z$ is equipped with the partially ordered net convergence structure of $(D)$-convergence. Suppose that, given $P_{0}\in\mathcal{P}(\Omega)$, for each $P\in\mathcal{P}(\Omega)$ such that $P\geq P_{0}$, we may find a common (unique) regular for the convergence of the series $\{\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)\}_{\alpha\in P}$ (for every choice function $\delta$), which is assumed convergent for each $P$ chosen this way. Then, $f$ is Sion integrable.
Proof.
By definition, as $f$ is $S^{*}$-integrable on $\Omega$, there exists a $(D)$-sequence $\{a_{ij}\}_{ij}$ in $Z$ such that, for each $\varphi\in\mathbb{N}^{\mathbb{N}}$ we may choose $P_{0}\in\mathcal{P}(\Omega)$ such that, for each $P\geq P_{0}$ in $\mathcal{P}(\Omega)$,
$|\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)-z|\leq\bigvee_{i=1}^{\infty}a_{i\varphi(i)},$
where $z\in Z$ is the $S^{*}$-integral of $f$ on $\Omega$.
Choose a common $(D)$-sequence for the convergence of the nets $\{\sum_{j\in J,\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})\}_{J\in\mathcal{F}(\mathbb{N})}$. That is, there exists, for all $P\in\mathcal{P}(\Omega)$, a $(D)$-sequence $\{b_{ij}\}_{ij}$ in $Z$ such that, for each $\varphi\in\mathbb{N}^{\mathbb{N}}$ we may choose $J=J(P)\in\mathcal{F}(\mathbb{N})$ such that, for each $J^{{}^{\prime}}\in\mathcal{F}(\mathbb{N})$ which contains $J$,
$|\sum_{j\in J^{{}^{\prime}},\alpha\in P}f(\delta_{\alpha_{j}})\mu(\alpha_{j})-\sum_{\alpha\in P}f(\delta_{\alpha})\mu(\alpha)|\leq\bigvee_{i=1}^{\infty}b_{i\varphi(i)}.$ (1)
By Fremlin lemma (in this setting, the Dedekind completeness and weak-$\sigma$-distributiveness are not needed), there exists $\{c_{ij}\}_{ij}$ a $(D)$-sequence in $Z$ such that:
$\bigvee_{i=1}^{\infty}a_{i\varphi(i)}+\bigvee_{i=1}^{\infty}b_{i\varphi(i)}\leq\bigvee_{i=1}^{\infty}c_{i\varphi(i)}.$ (2)
Now, fix $P_{0}^{(Si)}=P_{0}$ and define a truncation $\Delta_{0}$ by:
$\Delta_{0}(Q)=\{\alpha_{j}\}_{j\in J(Q)},$
where $J(Q)\in\mathcal{F}(\mathbb{N})$ is given in the convergence statement above for $Q\geq P_{0}$. Therefore, if $(P,\Delta)>>(P_{0}^{(Si)},\Delta_{0})$, we have that, by equations 1 and 2,
$|x-\sum_{\alpha\in\Delta(P)}f(\delta_{\alpha})\mu(\alpha)|\leq\bigvee_{i=1}^{\infty}c_{i\varphi(i)},$
for each $\varphi\in\mathbb{N}^{\mathbb{N}}$ as above. That is, $f$ is Sion integrable on $\Omega$ and:
$\int_{\Omega}^{S^{*}}fd\mu=z=\int_{\Omega}^{(Si)}fd\mu$
An interesting open question is if we may prove the same without supposing the existence of the unique $(D)$-sequence for each series of refining partitions. Maybe a possibility would be a conjunction of $Z$ being super Dedekind complete with another type of "order separability" property, but we leave this as an open problem.
We finish this section now, and consider some convergence theorems for the integrals introduced so far, and justifications for them.
4 Convergence Theorems for Net Riemann and Related Integrals.
To obtain more refined properties of the net Riemann integral and correlated concepts introduced in previous sections, we need some convergence theorems, specially (order and topological) uniform ones. The purpose of these convergence theorems are twofold:
@@
    Obtaining uniform convergence theorems for the Net Riemann integral, and specially the $S^{*}$(-partition) integral allows us to compare them to topological and order based Lebesgue integrals in the literature, such as Rodríguez Salazar (1985), Fernández y Fernández-Arroyo (1992), Fernández y Fernández-Arroyo et al. (1995), Haluška and Hutník (2012), Pavlakos (1983), Wright (1969) and to other order/topological based and sequentially/net defined integrals in the literature, unifying an and extending a wide range of integration procedures;
  2. (ii)
    Based on the comparisons contained in item (i), we may obtain monotone and dominated convergence theorems and other limit type properties for these integrals, as well as initial steps for a general classification of such integration procedures.
To obtain these convergence theorems, we would like to make as few assumptions as possible on the integrability of the sequence of functions and (order/topological based) limits involved. For the order case, we will adopt a generalization of the uniform convergence theorem of Monteiro and Fernandez (2012), which is based on the proof of a related theorem in Riečan and Vrábelová (1998), to the Henstock-Kurzweil integral (with respect to the Lebesgue measure) which, compared to Riečan and Neubrunn (1997) - which extends the partially ordered operator-valued framework of Riečan and Vrábelová (1993) - posits less strong integrability conditions on the limit of the approximating functions. This theorem is given in the cited paper for the Lebesgue measure and the Kurzweil integral. We shown in this section that this theorem is actually valid for a wide class of net Riemann integrals, and even for a outside of null sets version of $S^{*}$-integral, which is necessary to compare it to the order-convergence integrals of Pavlakos (1983) and Pavlakos (1991) (originated in Pavlakos (1972)).
We then prove a general Henstock lemma and obtain from it monotone and dominated convergence theorems in the setting or Riesz spaces, and survey some topological versions of convergence theorems, such as the convergence theorems of Fleischer (1995), Fleischer (1993), which generalized Sion (1973) dominated convergence theorem, and the outside of null sets topological convergence theorem of Millington (1975), which we generalize for a specific class of partition refinement integrals of the net Riemann integral, which contains more integrals than the one exposed by Millington (which is a variant of the Sion integral of Sion (1973)). This last result in particularly important to obtain equivalences with Lebesgue integrals defined by a type of Egorov procedure of local uniform convergence. Finally, we prove a general convergence theorem in the setting of lattice normed spaces.
We now fix some conventions for this section (and subsections):